# Learning nonnegative matrix factorizations from compressed data

### v1

@@ 5.2. Synthetic data: effect of compression size
Figure 2. NMF recovery of the synthetic data with MU. MU on the uncompressed data achieve limiting similarity $1$ after a few but slow iterations (only the limiting level is shown as “Uncompressed”). Displays (c,d) show that MU on data-adapted and random two-sided sketched data also tend to the limiting similarity $1$. Across all methods, less compression (larger $k$) improves convergence.
5.3. Real-world data: performance comparison
[…]
@@ 5.3. Real-world data: performance comparison
We also compare our proposed methods with a “NMF with random projections” method proposed in [42] and in [36]. These works adapted the updates of [4] to the compressed NMF setting resulting in the updates:
${\mathbf{U}}$ $\leftarrow{\mathbf{U}}\circ\sqrt{\frac{({\mathbf{Y}}_{2}{\mathbf{A}}_{2}^{T}{\mathbf{V}})_{+}+({\mathbf{U}}{\mathbf{V}}^{T}{\mathbf{A}}_{2}{\mathbf{A}}_{2}^{T}{\mathbf{V}})_{-}}{({\mathbf{Y}}_{2}{\mathbf{A}}_{2}^{T}{\mathbf{V}})_{-}+({\mathbf{V}}^{T}{\mathbf{A}}_{2}{\mathbf{A}}_{2}^{T}{\mathbf{V}})_{+}}}$ (5.1)
@@
and similar for ${\mathbf{V}}$, see equations (3.28) and (3.29) of [42]. The work of [42] propose to use the updates (5.1) where ${\mathbf{A}}_{1}$ and ${\mathbf{A}}_{2}$ are chosen to be Gaussian matrices. In this case we denote these updates “WL” in the legends. The work of [36] also proposed using the randomized rangefinder procedure [14] as in our Corollary 3.6 to choose the matrices ${\mathbf{A}}_{1}$ and ${\mathbf{A}}_{2}$. In this case we denote this method “TS” in the legends. Observe that these iterations are approximately two times faster than those of MU.
@@
Figure 3. (a,c): recovery from random Gaussian measurements, averaged over $5$ runs; our two-sided MU methods lead to better convergence than WL [42]. (b,d): data-adapted methods with sketching matrices obtained with the randomized rangefinder algorithm (like in Corollary 3.6); our two-sided multiplicative updates perform slightly better than TS [36] after enough iterations.
@@
Since it might be infeasible to even run an uncompressed problem from a memory perspective, we do not specifically focus on time performance here. However, we note that while it typically requires less iterations for the uncompressed NMF to learn the factorization, the iterations themselves are considerably slower. In Figure 3 (c,d) it would take several hours to run the uncompressed experiment until $60,000$ iterations, while the other methods take at most several minutes, so we show only the first $10^{3}$ iterations for uncompressed NMF. For the Faces dataset (Figure 3 (b)), it took $7$ sec to run compressed MU and $6$ sec to run GD until $60,000$ iterations, and we can see that $10$ times less iterations would have been enough for approximately same quality. The uncompressed method took $8$ sec for the plotted $10^{3}$ iterations, so at least $8$ min would be required to run it for $60,000$ iterations.
5.4. Real-world data: interpretable low rank decompositions
@@
5.5. Real-world data: choosing regularization parameter $\lambda$
Figure 5. Effect of regularization parameter $\lambda$ on the MU algorithm (4.6). 20News dataset compressed with data-adapted one-sided measurements, $\sigma$ is chosen minimal so that we have ${\mathbf{A}}{\mathbf{A}}^{T}\geq-\sigma$. The absence of regularization compromises convergence and too strong regularization results in a higher loss.
@@ 6. Conclusion and future directions
There are multiple venues of future work stemming from our approach. For the two-sided measurements, we currently do not have a theorem that holds for the data matrices with approximately low nonnegative rank, like in the one-sided case. The experimental evidence clearly suggests similar more general results should hold in the two-sided case. Also, it would be interesting to explain theoretically why the two-sided compression is less sensitive to the regularization (in practice, we take $\lambda=0$ in the two-sided experiments, which significantly simplifies the updates from Corollary 4.3).
@@
Further, it is natural and meaningful to extend the framework to the compressed versions of high-order (tensor) problems. It has been recently shown [20, 39] that nonnegative tensor factorization result in more interpretable decomposition for naturally high-order data, such as temporal or multi-agent, than their matrix NMF counterparts. At the same time, the tensor methods are even more computationally demanding and would benefit from scalable approximations. Scalable versions of other NMF-based algorithms such as semi-supervised versions [1] or sparse NMF [20, 17] are also of interest.

### v2

@@ 5.2. Effect of compression size $k$ on recovery
Figure 2. NMF recovery of the synthetic data with MU. Displays (c,d) show that MU on data-adapted and random two-sided sketched data also tend to the limiting similarity $1.0$. Across all methods, less compression (larger $k$) improves convergence.
Specifically, we choose the compression matrix $\mathbf{A}$} to be (a) a random Gaussian matrix, (b) a random matrix with orthogonal rows, or (c) via the randomized SVD procedure (3). We employ the compressed MU algorithm as in Corollaries 4, 5, and 6 respectively. For the two-sided method (d), we choose both compression matrices to be random Gaussian matrices and $\lambda=0$. For the one-sided compression, we take $\lambda=0.1$. Here, we report cosine similarity loss. Synthetic data has exactly low nonnegative rank $k\leq 20$. For reference, MU on the uncompressed data achieve limiting similarity very close to $1$.
Our the theoretical results suggest the possibility of exact recovery from the compressed data in two cases: (1) if two-sided compression was employed (Theorem 1), or (2) if the compression matrix is such that $P_{{\mathbf{A}}}^{\perp}{\mathbf{X}}$ is zero, for example, if the compression matrix is learned via the randomized rangefinder algorithm (Corollary 1). Correspondingly, the one-sided data-adapted compression and two-sided random compression attain exact reconstruction at different rates depending on the amount of compression (Figure 2 (c,d)). Two-sided compression requires twice more memory for the same sketch size but works with generic (data-oblivious) measurements that can be advantageous for various reasons as they do not require access to data or another pass over the data.
For the most compact one-sided data-oblivious measurements, Theorems 4 and 2 — corresponding to the measurements of Figure 2 (a,b) respectively — imply that the additional error might appear as a result of compression (from the term with $\|P^{\perp}_{{\mathbf{A}}^{T}}{\mathbf{X}}\|^{2}_{F}$). Here, we show that on an “easy” synthetic problem oblivious one-sided measurements can be also used: the compressed MU algorithm results in a good, although not perfect, recovery (Figure 2 (a,b)). The amount of additional loss depends on the amount of compression: indeed, for larger $k$ the dimension of the row-space of $\mathbf{A}$} is larger and the operator $P^{\perp}_{{\mathbf{A}}^{T}}$ projects the data into the smaller space, leading to smaller additional losses in the bounds such as (7) and (13).
In summary, we observe that less compression leads to a faster or better convergence performance in all the scenarios of Figure 2.
5.3. Effect of regularization parameter $\lambda$ on recovery
Figure 3. Effect of regularization parameter $\lambda$ on the MU algorithm (27). 20News dataset compressed with data-adapted one-sided measurements, $\sigma$ is chosen minimal so that we have ${\mathbf{A}}{\mathbf{A}}^{T}\geq-\sigma$. The absence of regularization compromises convergence and too strong regularization results in a higher loss.
5.4. Recovery methods performance comparison on real-world data
@@
We compare our proposed methods with a “NMF with random projections” method proposed in [45] and in [39]. These works adapted the updates of [4] to the compressed NMF setting resulting in the updates:
(29) ${\mathbf{U}}$ $\leftarrow{\mathbf{U}}\circ\sqrt{\frac{({\mathbf{Y}}_{2}{\mathbf{A}}_{2}^{T}{\mathbf{V}})_{+}+({\mathbf{U}}{\mathbf{V}}^{T}{\mathbf{A}}_{2}{\mathbf{A}}_{2}^{T}{\mathbf{V}})_{-}}{({\mathbf{Y}}_{2}{\mathbf{A}}_{2}^{T}{\mathbf{V}})_{-}+({\mathbf{V}}^{T}{\mathbf{A}}_{2}{\mathbf{A}}_{2}^{T}{\mathbf{V}})_{+}}}$
@@
and similar for $\mathbf{V}$}, see equations (3.28) and (3.29) of [45]. The work of [45] propose to use the updates (29) where ${\mathbf{A}}_{1}$ and ${\mathbf{A}}_{2}$ are chosen to be Gaussian matrices. In this case we denote these updates (WL) in the legends. The work of [39] also proposed using the randomized rangefinder procedure [15] as in our Corollary 1 to choose the matrices ${\mathbf{A}}_{1}$ and ${\mathbf{A}}_{2}$. In this case we denote this method (TS) in the legends.
For our two-sided (MU) method we solve a simplified version of the two-sided objective (17) with $\lambda=0$. We take ${\mathbf{A}}_{1}$ and ${\mathbf{A}}_{2}$ to be random Gaussian matrices in the oblivious case or according to the randomized rangefinder procedure as in Corollary 1 in the adaptive case. For the data-adapted one-sided (MU) method, we take $\lambda=0.1$ and solve the problem (26). Then, we include the recovery via projected gradient descent (GD), as described in Section 4.3 with a step size of $\alpha=.001$.
@@
Figure 4. (a,c): recovery from random Gaussian measurements, averaged over 55 runs; our two-sided MU methods lead to better convergence than WL [45]. (b,d): data-adapted methods with sketching matrices obtained with the randomized rangefinder algorithm (like in Corollary 1); our two-sided MU perform slightly better than TS [39] after enough iterations.
@@
Since it might be infeasible to even run an uncompressed problem from a memory perspective, we do not specifically focus on time performance here. However, we note that while it typically requires less iterations for the uncompressed NMF to learn the factorization, the iterations themselves are considerably slower. In Figure 4 (c,d) it would take several hours to run the uncompressed experiment until $60,000$ iterations, while the other methods take at most several minutes, so we show only the first $10^{3}$ iterations for uncompressed NMF. For the Faces dataset, it took $16$ sec to run compressed MU and $12$ sec to run GD until $60,000$ iterations, and we can see that $10$ times less iterations would have been enough for approximately same quality. The uncompressed method took $7$ sec for the plotted $10^{3}$ iterations, so at least $7$ min would be required to run it for $60,000$ iterations.
See Appendix A for additional information about the expreriments from this section, including the time and resulting cosine similarity performance of each method.
5.5. Recovery methods produce interpretable decompositions on real-world data
@@ 6. Conclusion and future directions
There are multiple venues of future work stemming from our approach. For the two-sided measurements, we currently do not have a theorem that holds for the data matrices with approximately low nonnegative rank, like in the one-sided case. The experimental evidence clearly suggests similar more general results should hold in the two-sided case. Also, it would be interesting to explain theoretically why the two-sided compression is less sensitive to the regularization (in practice, we take $\lambda=0$ in the two-sided experiments, which significantly simplifies the updates from Corollary 4).
An important challenge is to meaningfully transfer the obtained theoretical connection between the optimal values of the standard and compressed NMF to the decompositions that we learn in practice. The obstacle is in poor guarantees for the convergence of the NMF learning methods. Recently, in [28], it was shown that a regularized and component-wise bounded version of MU indeed converges to a first-order optimal point with a controlled number of iterations. It would be reasonable to consider an analogue of this version of MU for the compressed objective and tighten the algorithmic connection between the full and compressed problems through it.
@@
Further, it is natural and meaningful to extend the framework to the compressed versions of high-order (tensor) problems. It has been recently shown [21, 42] that nonnegative tensor factorization result in more interpretable decomposition for naturally high-order data, such as temporal or multi-agent, than their matrix NMF counterparts. At the same time, the tensor methods are even more computationally demanding and would benefit from scalable approximations. Scalable versions of other NMF-based algorithms such as semi-supervised versions [1] or sparse NMF [21, 18] are also of interest. Finally, while in this work we focus solely on linear sketching techniques, it is interesting to investigate both theoretical and empirical properties of nonlinear dimension reduction for the NMF and nonnegative tensor factorization problems.
Appendix A Additional numerical evidence: time comparison
Here, we include additional summarizing information for the experiments from Section 5.4.
Method Oblivious/
Adapted
Memory needed Time per
iteration
Similarity
after $10^{3}$
iterations
Final
similarity
Uncompressed 100% 6.79 .9872 N/A
WL O 10% 0.26 .8279 .8280
TS A 10% 0.27 .9439 .9863
MU 2-sided O 10% 0.23 .9322 .9631
MU 1-sided A 5% 0.27 .9626 .9848
MU 2-sided A 10% 0.23 .9347 .9865
GD A 5% 0.2 .9806 .9849
Table 2. Summary of results for the Faces dataset. Oblivious measurements (O) advantageous as they can be applied to any data, data-adapted (A) typically result in better performance. All oblivious methods were run until $10^{6}$ iterations total and all data-adapted methods until $6*10^{4}$ iterations total; final similarity is reported at the last iteration. Time is reported in microseconds. Similarity above $98\%$ is highlighted.
Method Oblivious/
Adapted
Memory needed Time per
iteration
Similarity
after $10^{3}$
iterations
Final
similarity
Uncompressed 100% 50.6 .2018 N/A
WL O 4% 0.73 .0024 .0024
TS A 4% 0.73 .1648 .1655
MU 2-sided O 4% 1.5 .0898 .1013
MU 1-sided A 2% 1.84 .0935 .1471
MU 2-sided A 4% 1.5 .1167 .1709
Table 3. Summary of results for the 20News dataset. All compressive methods were run until $6*10^{4}$ iterations total; final similarity is reported at the last iteration. Time is reported in milliseconds. Similarity above $17\%$ is highlighted (it is a hard dataset, but this similarity can be still enough for topic modeling tasks).